OpenStudy (zasharra):

solve for x: 5/x^2-4+2/x=2/x-2

OpenStudy (whpalmer4):

Okay, we can't see that file. Does the problem look like this? \[\frac{ 5}{x^2-4}+\frac{2}{x}=\frac{2}{x-2}\]If so, when you type it all on a single line, you need to put parentheses around the numerators and denominators: 5/(x^2-4) + 2/x = 2/(x-2) By the order of operator precedence, what you wrote actually means \[ \frac{5}{x^2}-4+\frac{2}{x}=\frac{2}{x}-2\]which is entirely different, I hope you'll agree!

OpenStudy (zasharra):

@whpalmer4 it's the first way you typed it.

OpenStudy (zasharra):

@whpalmer4 yes it's the first one

OpenStudy (zasharra):

is it 8?

OpenStudy (zasharra):

nvm!!! im right!!

OpenStudy (whpalmer4):

Let's check it and find out! \[\frac{ 5}{x^2-4}+\frac{2}{x}=\frac{2}{x-2}\]substitute \(x = 8\) \[\frac{ 5}{8^2-4}+\frac{2}{8}=\frac{2}{8-2}\]\[\frac{5}{60}+\frac{2}{8}=\frac{2}{6}\]Looks like 240 is a common denominator\[\frac{4}{4}*\frac{5}{60}+\frac{30}{30}*\frac{2}{8} = \frac{40}{40}*\frac{2}6\]\[\frac{20+60}{240}=\frac{80}{240}\checkmark\]Yes, \(x=8\) is the solution